In triangle RST, RS= 2x+10, ST = 3x-2 and RT = (1/2)x +28. if triangle is equiangular, what is the value of x ?
A 2 B 5 1/3 C 6 D 12 E 34

Answer by nyc_function(2741)

If the triangle is equiangular, then all sides must be the same length.
Equate any two sides that you like and solve for x.
I will equate 2x + 10 to 3x - 2.
2x + 10 = 3x - 2
2x - 3x = -2 - 10
-x = -12
x = -12/-1
x = 12
CHOICE D is correct.